# Machine Unlearning in Speech Emotion Recognition via Forget Set Alone

### v1

@@ 1 Introduction
Human speech is an information-rich resource that can provide valuable insights into paralinguistic cues such as emotional states. The growing trend towards more naturalistic human-machine interactions has made the ability to automatically understand and interpret emotions from speech more relevant than ever [1]. Speech Emotion Recognition (SER) has been proposed to automatically identify emotional states from human speech using machine learning methods [2]. SER is promising for manifold applications, such as car-driving [3], education [4], healthcare [5], etc. More recently, end-to-end models, e. g., Wav2Vec [6] and HuBERT [7], trained with self-supervised learning on large-scale datasets have demonstrated a strong capability in extracting abstract emotion-relevant representations. Therefore, they can yield superior performance for SER when fine-tuned on emotional speech datasets. Such good performance promotes the applications of SER based on streaming speech data from various devices, including online platforms, wearable devices, and many others.
@@
We propose applying a machine unlearning approach using only the data to be forgotten for SER. The proposed approach can (i) train a model to forget the data to be erased, and (ii) maintain model performance using generated adversarial samples that capture the emotional class characteristics of the remain set. The experimental results demonstrate that the proposed approach can train an SER model to forget the erased data, resulting in a reduction of SER accuracy to approx. $0.0\%$ for erased data, while still achieving adequate performance on the original test set.
@@
For SER, the study in [12] proposed a weight averaging method to combine multiple models, each of which is trained on a data shard. Differently, we focus on instance unlearning of one SER model with the data to be forgotten only, avoiding potential privacy risk of using the remain set. Inspired by the study for image classification [16], we maintain the model performance for SER using generated adversarial samples and elastic model weights.
@@ 2.1 Misclassification of Forget Set
To misclassify $D_{e}$, an effective way is to randomly relabel the data with wrong labels that are different from their original true ones [15]. In this regard, we randomly generate wrong labels for data samples in $D_{e}$. This random relabelling procedure leads to $\hat{D}_{e}:\{\bm{x_{e}},\hat{y}_{e}\}$, $\hat{y}^{i}_{e}\neq y^{i}_{e}$, where $i$ means the sample index. The model is then trained on $\hat{D}_{e}$ for misclassification using the loss function:
$\vskip-2.84526pt\mathcal{L_{\mbox{mis}}}=\mathcal{L_{\mbox{CE}}}(f(\bm{x_{e}}),\hat{y}_{e}),$ (1)
where $\mathcal{L_{\mbox{CE}}}$ denotes the cross-entropy loss.
@@ 2.2.3 Model Training
$\vskip-2.84526pt\mathcal{L}=\lambda_{1}\mathcal{L}_{\mbox{mis}}+\lambda_{2}\mathcal{L}_{\mbox{adv}}+\lambda_{3}\mathcal{L}_{\mbox{ewc}},$ (5)
@@ 3.2 Experimental Setup
Table 1: Model performance (UAR) on the forget set ($D_{e}$) and the validation/test set when training models on the train(ing) / train(ing)+val(idation) sets. The NN denotes the number of speech samples to be forgotten. Apart from fine-tuning a Wav2Vec 2.0 model, the baselines of remaining-data-involved unlearning (Remain. Unl.) and random labelling (Ran. Lab.) are compared with the proposed approaches using adversarial attacks (Adv.) and adversarial attacks + elastic model weights (Adv. + Ela.). The best performance of the proposed approach and Ran. Lab. are compared with significant tests (*: $p<0.001$ in a one-tailed z-test).
@@
Baselines. In machine unlearning, the training epoch is $15$ and the optimiser setting is the same as that in fine-tuning. The adversarial-attack-based machine unlearning approach is compared to two baselines of machine unlearning. (i) Remaining-data-involved unlearning. The model is trained not only with random labelling, but also on the remain set $D_{r}$, thereby the loss function is $\mathcal{L}_{\mbox{mis}}+\mathcal{L}_{\mbox{CE}}(f(\bm{x_{r}},y_{r}))$ [15]. (ii) Random labelling. The model is trained with the forget set and randomly generated wrong labels using (1[15].
@@ 3.4 Results Comparison
Compared to the baselines, the proposed two approaches using adversarial attacks can make the model forget the knowledge learnt from $D_{e}$. Both approaches perform on $D_{e}$ with UARs not higher than $0.110$. Both approaches also outperform the random labelling method, indicating the effectiveness of adversarial attacks in augmenting the data and simulating the data distribution of the remain set. When comparing the two proposed approaches, using elastic model weights can further improve the model performance. The reason can be that the EWC method regulates the models to only update unimportant model parameters. Finally, the models in the proposed two approaches perform worse when $N$ increases, which is reasonable. We can still see significant improvement of the model performance compared to random labelling when $N=10,30,50$ ($p<0.001$ in a one-tailed z-test).
@@ 4 Conclusion and Future Work
This work proposed a machine unlearning approach using adversarial attacks to protect data privacy hidden in emotional speech. The proposed approach utilises the forget set only, and generates adversarial samples to help the model preserve the knowledge learnt from the remain set. The weights of the speech emotion recognition model are updated by considering parameter importance during unlearning, thereby preserving more knowledge of the remain set. The experimental results indicate that the proposed approach can effectively train the model to forget the date to be erased and still perform well on unseen speakers’ data for emotion recognition. In future efforts, we will validate the effectiveness of the approach in various speech emotion recognition models. We will also investigate the approach for forgetting specific emotional classes and sensitive information, such as gender.

### v2

@@ 1 Introduction
The growing trend towards more naturalistic human-machine interactions has made the ability to automatically understand and interpret emotions from speech more relevant than ever [1]. Speech Emotion Recognition (SER) has been proposed to automatically identify emotional states from human speech using machine learning methods [2], and it has proven useful in applications such as healthcare, education, etc. More recently, end-to-end models, e. g., Wav2Vec [3] and HuBERT [4], trained with self-supervised learning on large-scale datasets have demonstrated superior performance for SER when fine-tuned on emotional speech datasets. Such good performance promotes the applications of SER based on streaming speech data from various devices, including online platforms, wearable devices, and many others.
@@
We propose applying a machine unlearning approach using only the data to be forgotten for SER. The proposed approach can (i) train a model to forget the data to be erased (a reduction of accuracy to approx. $0.0\%$ for erased data), and (ii) maintain model performance using generated adversarial samples that augment the training data.
@@
For SER, the studies in [9, 13] proposed a weight averaging method to combine multiple models, each of which is trained on a data shard. Differently, we focus on instance unlearning of one SER model with the data to be forgotten only, avoiding potential privacy risk of using the remain set. Inspired by the study for image classification [14], we maintain the model performance for SER using generated adversarial samples and elastic model weights.
@@ 2.1 Misclassification of Forget Set
To misclassify $D_{e}$, an effective way is to randomly relabel the data with wrong labels that are different from their original true ones [12]. In this regard, we randomly generate wrong labels for data samples in $D_{e}$. This random relabelling procedure leads to $\hat{D}_{e}:\{\bm{x_{e}},\hat{y}_{e}\}$, $\hat{y}^{i}_{e}\neq y^{i}_{e}$, where $i$ means the sample index. The model is then trained on $\hat{D}_{e}$ for misclassification using the cross-entropy loss function $\mathcal{L_{\mbox{CE}}}$:
$\vskip-2.84526pt\mathcal{L_{\mbox{mis}}}=\mathcal{L_{\mbox{CE}}}(f(\bm{x_{e}}),\hat{y}_{e}).$ (1)
@@ 2.2.3 Model Training
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\mbox{mis}}+\lambda_{2}\mathcal{L}_{\mbox{adv}}+\lambda_{3}\mathcal{L}_{\mbox{ewc}},\vskip-2.84526pt$ (5)
@@ 3.2 Experimental Setup
Table 1: Model performance (UAR) on $D_{e}$ and the validation/test set. NN denotes the number of samples to be forgotten. Apart from fine-tuning Wav2Vec 2.0, remaining-data-involved unlearning (Remain. Unl.) and random labelling (Ran. Lab.) are compared with the proposed approaches using adversarial attacks (Adv.) and adversarial attacks + elastic model weights (Adv. + Ela.). The best performance of the proposed approaches and Ran. Lab. are compared with significant tests (*: $p<0.001$ in a one-tailed z-test).
@@
Baselines. In machine unlearning, the training epoch is $15$ and the optimiser setting is the same as that in fine-tuning. WavLM with adversarial attacks and elastic model weights was trained for 20 epochs to erase the forget set. The adversarial-attack-based machine unlearning approach is compared to two baselines. (i) Remaining-data-involved unlearning. The model is trained not only with random labelling, but also on $D_{r}$, thereby the loss function is $\mathcal{L}_{\mbox{mis}}+\mathcal{L}_{\mbox{CE}}(f(\bm{x_{r}},y_{r}))$ [12]. (ii) Random labelling. The model is trained with the forget set and randomly generated wrong labels using (1[12].
@@ 3.4 Results Comparison
Compared to the baselines, the proposed two approaches using adversarial attacks can make the model forget the knowledge learnt from $D_{e}$. Both approaches perform on $D_{e}$ with UARs not higher than $0.110$. Both approaches also outperform the random labelling method, indicating the effectiveness of adversarial attacks in augmenting the training data. When comparing the two proposed approaches, using elastic model weights can further improve the model performance. The reason can be that the EWC method regulates the models to only update unimportant model parameters. Finally, the models in the proposed two approaches perform worse when $N$ increases, which is reasonable. We can still see significant improvement of the model performance compared to random labelling when $N=10,30,50$ ($p<0.001$ in a one-tailed z-test).
Finally, we compare the approaches across Wav2Vec 2.0, HuBERT [4], and WavLM [22] in Table 2. The proposed approaches using adversarial attacks w/ and wo/ elastic model weights mostly outperform the other two methods, including remain-data-involved unlearning and random labelling. This further indicates the effectiveness of the proposal approaches.
Table 2: Model comparison on the test set of DEMoS. Number of speech samples to be forgotten is 30. (*: $p<0.001$ in a one-tailed z-test)
Wav2Vec 2.0 HuBERT WavLM
Method $D_{e}$ Test $D_{e}$ Test $D_{e}$ Test
Finetune 0.897 0.909 0.908
Remain Unl. 0.486 0.910 0.060 0.771 0.071 0.810
Ran. Lab. 0.071 0.506 0.286 0.629 0.179 0.599
Adv. 0.000 0.747* 0.000 0.473 0.167 0.723*
Adv. Ela. 0.000 0.731 0.107 0.740* 0.143 0.696
@@ 4 Conclusion and Future Work
This work proposed a machine unlearning approach using adversarial attacks to protect data privacy hidden in emotional speech. The proposed approach utilises the forget set only, and generates adversarial samples to help the model on data augmentation. The weights of the speech emotion recognition model are updated by considering parameter importance during unlearning, preserving more knowledge of the remain set. The experimental results indicate that the proposed approach can effectively train the model to forget the date to be erased and still perform well on unseen speakers’ data for emotion recognition. The performance improvement is mainly from the data augmentation effect of adversarial samples, as confirmed by our experiments. We observe that simply increasing the size of the forgotten set degrades performance, indicating that more principled adversarial generation strategies (e. g., generation of adversarial data from representative prototypes) are needed. As the forget set is randomly selected from a single speaker, we will explore the speaker variability in future work. Finally, we will investigate the approach for forgetting specific emotional classes and sensitive information, such as gender.